mcSqueaks:

After 2 hours, a car travels 70 miles. After 2.25 hours in the same trip, the car travels 78.75 miles. I think I need point slope form and slope intercept form.

mcSqueaks:

I have the formulas

mcSqueaks:

Slope intercept = y=m(x+b) and point slope form y − y1 = m(x − x1)

QuestionCoveBot:

So it traveled 70 miles in 2 hours. It traveled 35 miles per hour. So after .25 hours (15 minutes), they traveled an extra 8.75 miles.

mcSqueaks:

Sorry so what is the equation

1 attachment
mcSqueaks:

Ignore the highlighted plz

QuestionCoveBot:

Are you trying to find the slope?

mcSqueaks:

I dunno that's all it says is what I sent you

QuestionCoveBot:

It's a positive slope, so it will be something with a positive sign.

QuestionCoveBot:

And something with 35 and 70.

QuestionCoveBot:

And for your information, you clicked the right answer.

mcSqueaks:

Got it thanks!

QuestionCoveBot:

Mhm, you're welcome.